A-6nC particle is moving at 4.0 x 10⁴ m/s along the positive z-axis. It passes a region indeflected where there is uniform electric and magnetic fields. The electric field is 5.75 x 10⁵ N/C along the negative x-axis. What is the magnitude and direction of the magnetic field?

1
Expert's answer
2022-03-22T11:00:19-0400

The charged particle in uniform electric and magnetic field moves along straight line, so net force on charge is zero. Hence

"qE=qvB"

The magnitude of magnetic field

"B=\frac{E}{v}=\frac{5.75*10^5}{4.0*10^4}=14.4\;\rm T"

The magnetic field is directed toward negative "y"-axis.